OpenStudy (pinkamenadiianepie):

i need help solving an equation.... -7(k + 9) = 9 (k - 5) - 14k

OpenStudy (welshfella):

OK first you need to eliminate the parentheses using the distributive law can you do that?

OpenStudy (pinkamenadiianepie):

no

OpenStudy (welshfella):

OK you multiply what's in the parenthese by the number in front -7(k + 9) = 9(k - 5) - 14k -7*k - 7*+9 = 9*k + 9*-5 - 14k can you continue (perform the multiplications)

OpenStudy (pinkamenadiianepie):

w ( so do i divide them

OpenStudy (welshfella):

no you multiply the left side becomes -7k - 63 ( becuase -7*k = -7k and -7*9 = -63) now do the same process on the right side of the equation

OpenStudy (welshfella):

what is 9*k + 9 * -5?

OpenStudy (kailey7277):

First, you would use the distributive property. You would multiply -7 to everything in the first set of parentheses. \[-7*k=-7k\] \[-7*9=-63\] Put them together and you have: \[-7k-63\] Then, you would do the same thing to the parentheses on the other side. \[9*k=9k\] \[9*(-5)=-45\] put that together and you have 9k-45. Don't forget about the -14k, though. So now you would have: \[-7k-63=9k-45-14k\] Now, you need to combine like terms on the right side. \[9k-14k=-5k\] Then, you will have: \[-7k-63=-5k-45\] Now, you have to add 7k to both sides. \[-7k-63=-5k-45\] \[+7k.........+7k\] \[-63=2k-45\] Next, you would add 45 to both sides to get the variable alone. \[-63=2k-45\] \[+45.........+45\] \[-18=2k\] You would then divide both sides by 2. \[-9=k\] You should probably flip the equation so that the variable i on the left side: \[k=-9\]